OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x - y = -4 4x - 3y = -6

OpenStudy (anonymous):

4x - 3(-12x - 4) = -6 4x - 3(12x + 4) = -6 4(12x + 4) - 3y = -6 4(-12x - 4) - 3y = -6

OpenStudy (cakt07):

@chelliegrace help O_O